Problem 18-09

Shamrock Construction Company has entered into a contract beginning January 1, 2020, to build a parking complex. It has been estimated that the complex will cost $601,000 and will take 3 years to construct. The complex will be billed to the purchasing company at $903,000. The following data pertain to the construction period.

2020

2021

2022

Costs to date $270,450 $492,820 $613,000
Estimated costs to complete 330,550 108,180 –0–
Progress billings to date 273,000 550,000 903,000
Cash collected to date 243,000 500,000 903,000


(a) Using the percentage-of-completion method, compute the estimated gross profit that would be recognized during each year of the construction period. (If answer is 0, please enter 0. Do not leave any fields blank.)

Gross profit recognized in 2020 $

Gross profit recognized in 2021 $

Gross profit recognized in 2022 $


(b) Using the completed-contract method, compute the estimated gross profit that would be recognized during each year of the construction period. (If answer is 0, please enter 0. Do not leave any fields blank.)

Gross profit recognized in 2020 $

Gross profit recognized in 2021 $

Gross profit recognized in 2022 $
Answer #1

Requirement a

Gross profit recognized in 2020 $      135,900
Gross profit recognized in 2021 $      111,740
Gross profit recognized in 2022 $       42,360

Requirement b

Gross profit recognized in 2020 $ 0.00
Gross profit recognized in 2021 $0.00
Gross profit recognized in 2022 $ 290,000*

*Total Revenue - Total cost [903000-613000]

Working for Percentage of completion method

Working 2020 2021 2022
A Contract Price $ 903,000 $ 903,000 $ 903,000
B Cost Incurred to Date $ 270,450 $ 492,820 $ 613,000
C Estimated cost yet to be incurred to complete the contract $ 330,550 $ 108,180 $ 0
D = B+C Total Cost $ 601,000 $ 601,000 $ 613,000
E = (D/B) x 100 % of Completion 45% 82% 100%
F = A x E Revenue to date $ 406,350 $ 740,460 $ 903,000
G Revenue of Previous year $ 0 $ 406,350 $ 740,460
H = F - G Net Revenue this year $ 406,350 $ 334,110 $ 162,540
I Cost to date $ 270,450 $ 492,820 $ 613,000
J Cost to date of previous year $ 0 $ 270,450 $ 492,820
K = I - J Net Cost for the year $ 270,450 $ 222,370 $ 120,180
L = H - K Gross Profits $ 135,900 $ 111,740 $ 42,360